An oriental rug is 52 yrs old an Persian rug is 16 yes old. How many years ago was the oriental rug four times as old as the Persian rug.

Answer by ewatrrr(24785)

52-x = 4(16-x)
3x = 64-52
3x = 12
x = 4yr ago
Checking:
4 years ag0...
48 = 4(12) CHECKS!